# SyncLipMAE: Contrastive Masked Pretraining for Audio–Visual Talking-Face Representations

### v1

@@ 1 Introduction
Talking-face video underpins human–computer interaction, accessibility, content creation, and post-production. Practical systems must jointly reason about linguistic content, audio–visual timing, video editing to match a target speech track or reference motion, and facial expressions/actions. Classic synchronisation pipelines such as SyncNet and subsequent Transformer-based variants (e.g., VocaLiST) focus on detecting in-/out-of-sync pairs or removing stream lags, while recent audio–visual pretraining methods (e.g., AV-HuBERT) have shown strong transfer to lip reading and AV-ASR. Yet, these advances do not directly yield a token-level, stream-synchronised representation that cleanly separates identity, speech-synchronised mouth motion, and other facial dynamics—capabilities that are essential for both analysis and controllable generation.
Most prior work tackles one facet at a time. Visual or video masked modeling (MAE/VideoMAE) excels at reconstruction but is modality-specific and does not enforce audio–visual alignment; contrastive AV learning improves correspondence but is typically framed as binary sync classification rather than a shared token space; and dubbing systems (e.g., Wav2Lip, LatentSync, MuseTalk) optimise for generation quality and lip accuracy but are not designed to provide a unified interface that can be driven interchangeably by audio or reference motion within a single model. As a result, existing solutions often rely on wider temporal windows or task-specific heads, and they lack a factorised, alignment-aware representation that transfers seamlessly across synchronisation, understanding, and dubbing.
@@
Beyond analysis, SyncLipMAE enables unified visual dubbing: using a WanVACE-style DiT backbone for masked video inpainting, we inject either aligned audio tokens or vocal-motion tokens after patch embedding via a lightweight AudioPack, so the same model admits no-difference audio- or video-driven control. This preserves identity and pose while re-synchronising the mouth region, and it naturally benefits from the shared token space learned during pretraining.
@@ 2 Related Work
A long line of work formulates A/V learning as correspondence or temporal alignment.  Arandjelovic & Zisserman (2017) learns generic audio–visual correspondence from unlabelled video, while  Morgado et al. (2020) casts synchronisation as in-time vs. out-of-time discrimination within the same clip. For faces,  Chung & Zisserman (2016) introduces a two-stream embedding for lip–audio alignment and lag estimation; recent transformers such as  Kadandale et al. (2022) improve robustness across speech/singing with longer-range context.
@@
Contrastive pretraining aligns paired modalities at scale (e.g., CLIP  Radford et al. (2021)), while masked autoencoders (MAE He et al. (2022)/VideoMAE Tong et al. (2022)) show that reconstructing heavily masked inputs yields strong visual/video representations. These paradigms are complementary and now standard building blocks for multimodal pretraining.
@@
Closer to our setting, MARLIN Cai et al. (2023) applies masked autoencoding to facial videos, using facial-region-guided masking to learn a universal face encoder transferable to expression recognition, deepfake detection, etc. Subsequent variants adapt MAE-style pretraining to dynamic facial expression recognition under limited labels Sun et al. (2023). While these works focus on reconstruction-only objectives within the facial domain, we couple MAE with audio–visual contrast to align speech-driven dynamics, and structure the representation to separate identity and motion factors.
@@
From early end-to-end lipreading (LipNet) to large-scale self-supervised A/V pretraining (AV-HuBERT) and automated VSR/AVSR recipe design (Auto-AVSR Ma et al. (2023)), results consistently show that robust visual streams improve recognition and benefit from pretraining on unlabelled A/V data.
@@
Audio-driven talking-head synthesis has progressed from GAN-based pipelines with explicit sync critics (e.g., Wav2Lip Prajwal et al. (2020)) to diffusion in latent/image space (e.g., LatentSync Li et al. (2024)) and efficient diffusion heads (MuseTalk Zhang et al. (2024)). In parallel, foundation video generators and unified conditional DiT Peebles & Xie (2023) frameworks (WAN Wan et al. (2025); VACE Jiang et al. (2025)) provide scalable backbones and conditioning interfaces (Video Condition Unit, context adapters) for editing or T2V, which our dubbing setup leverages.
@@ 3.1 Overview and Design Goals
SyncLipMAE aims to learn synchronization-aware talking-face representations from unlabeled audio–visual streams. As illustrated in figure 1(a), our approach couples MAE-style masked image/video modeling on the visual stream with a CLIP-style cross-modal contrastive alignment that treats time-aligned vocal-motion and audio tokens as positives and temporally misaligned pairs as negatives, driving both modalities into a shared embedding space and yielding token-level audio–visual stream synchronization (§3.3). Per frame, a ViT-based encoder exposes three prompt tokens that explicitly encode the essential factors of a talking-face frame—identity $\mathbf{z}^{\mathrm{id}}$, vocal motion $\mathbf{z}^{\mathrm{voc}}$, and ambient motion $\mathbf{z}^{\mathrm{amb}}$—and the decoder reconstructs frames via cross-attention conditioned on these prompts.
@@
Let $\mathbf{x}_{1:T}\in\mathbb{R}^{T\times C\times H\times W}$ be video frames and $\mathbf{a}\in\mathbb{R}^{N}$ mono audio at sampling rate $f_{s}$. Each frame is patchified into $N_{p}$ patches and embedded to $D$-dimensional tokens. For MAE-style masking, denote by $\mathcal{V}_{t}\subset\{1,\dots,N_{p}\}$ the indices of visible patches at time $t$ and by $\mathcal{M}_{t}$ the masked ones, with $|\mathcal{V}_{t}|=N_{\mathrm{vis}}$ and $\mathcal{V}_{t}\cup\mathcal{M}_{t}=\{1,\dots,N_{p}\}$. The per-frame visible patch tokens are $\mathbf{P}^{\mathrm{vis}}_{t}\in\mathbb{R}^{N_{\mathrm{vis}}\times D}$. Let $\mathbf{e}^{\mathrm{mask}}\in\mathbb{R}^{D}$ be a learnable mask token; a restoration operator places $\mathbf{P}^{\mathrm{vis}}_{t}$ and copies of $\mathbf{e}^{\mathrm{mask}}$ at indices $\mathcal{M}_{t}$ to form a length-$N_{p}$ sequence for decoding. Per frame, the encoder also outputs three prompt tokens, $\mathbf{z}^{\mathrm{id}}_{t},\ \mathbf{z}^{\mathrm{amb}}_{t},\ \mathbf{z}^{\mathrm{voc}}_{t}\in\mathbb{R}^{D}$. For audio, we employ a pretrained speech encoder with $L$ layers that is kept frozen during pretraining; let $\{\mathbf{H}^{(\ell)}\}_{\ell=1}^{L}$ denote the hidden-state sequences from all layers. To obtain frame-wise alignment with the video, we resample the input waveform so that the encoder emits $T$ tokens per utterance (one token per video frame) at every layer, yielding $\mathbf{A}^{(\ell)}=\{\mathbf{a}^{(\ell)}_{t}\}_{t=1}^{T}$.
@@ 4 Downstream Adaptation
Method VFHQ Hallo3 $\mathrm{Sync}_{\mathrm{conf}}\,\uparrow$ FID $\downarrow$ FVD $\downarrow$ $\mathrm{Sync}_{\mathrm{conf}}\,\uparrow$ FID $\downarrow$ FVD $\downarrow$ MuseTalk Zhang et al. (2024) 0.6820 24.44 175.16 0.6904 13.80 86.46 LatentSync-1.5 Li et al. (2024) 0.7681 17.92 131.04 0.7725 10.59 78.89 A-SyncLipMAE (ours) 0.7830 16.46 100.44 0.8027 10.04 74.01 V-SyncLipMAE (ours) 17.22 109.91 10.82 77.57
We apply SyncLipMAE to four tasks: (i) audio–visual stream synchronization — infer frame-level lag and synchrony by comparing aligned audio tokens $\mathbf{A}_{t}$ with vocal-motion tokens $\mathbf{z}^{\mathrm{voc}}_{t}$; (ii) facial understanding — infer emotion and head/face actions from the vocal- and ambient-motion tokens; (iii) visual speech recognition (VSR) — infer text from the vocal-motion tokens; (iv) visual dubbing — edit a source face video to match target speech or a reference motion, supporting audio- or video-driven control within a single model.
@@ 4.2 Facial understanding: emotion and action
We use SyncLipMAE ’s motion factors per frame and keep the prediction head minimal. For each frame $t$, we form a motion descriptor by averaging the two factors, $\mathbf{u}_{t}=\tfrac{1}{2}\big(\mathbf{z}^{\mathrm{voc}}_{t}+\mathbf{z}^{\mathrm{amb}}_{t}\big)$, and apply a single linear classifier to obtain frame-level logits. Video-level logits are the mean of frame-level logits over the sampled frames; the final prediction is the softmax over the aggregated logits. This setup probes whether the representation cleanly separates speech-synchronized orofacial motion from audio-agnostic dynamics (e.g., blinks, head pose) while keeping the adaptation lightweight; datasets such as CelebV-HQ provide the emotion/action labels used in this setting.
@@ 4.3 Lip reading (VSR)
We use a visual-only sequence-to-sequence head on top of SyncLipMAE ’s vocal-motion tokens. Let $\mathbf{V}=\{\mathbf{z}^{\mathrm{voc}}_{t}\}_{t=1}^{T^{\prime}}$ be the input to a Conformer encoder, yielding hidden states $\mathbf{H}=\{\mathbf{h}_{t}\}_{t=1}^{T^{\prime}}$. A Transformer decoder autoregressively predicts subword units with cross-attention to $\mathbf{H}$.
@@
The total objective is the standard hybrid form
@@ 4.4 Video dubbing
We cast dubbing as masked video inpainting on face crops within a VACE-style unified interface: task inputs (reference frames and a spatiotemporal mouth mask) are organized as a Video Condition Unit and processed by a Diffusion-Transformer backbone with Concept Decoupling and Context Adapters, following VACE. Concretely, we instantiate the pipeline on WanVACE and keep its architecture, losses, and training schedule unchanged. WanVACE natively accepts multiple reference images (1–3 in the public release), exposed as a comma-separated list, which we leverage for identity control.
Control injection. Inspired by audio-conditioned injection in recent avatar generation, we adopt a minimal AudioPack placed after the VACE patch-embedding stage and before the first DiT VACE block. Per-frame control tokens—either audio tokens $\mathbf{A}_{t}$ produced by SyncLipMAE or reference vocal-motion tokens $\mathbf{z}^{\mathrm{voc}}_{t}$—are temporally encoded and projected to the DiT width, then added to the patch-embedded video token stream. No other WanVACE components are modified; switching $\mathbf{A}_{t}$ vs. $\mathbf{z}^{\mathrm{voc}}_{t}$ toggles audio-driven vs. video-driven control within the same model.
Long-form continuity via dual references. To support arbitrarily long, coherent dubbing, we provide two references: (i) an identity image (stable appearance prior), and (ii) the last frame of the previous segment (temporal bridge). During training we randomly drop the continuity reference with 10% probability, so the model can also initialize the first segment without a history frame.
Masking strategy. Following common lip-sync inpainting practice, we detect the facial region (chin to hairline) and mask the lower half as the editable area; to accommodate jaw excursions (post-edit lip motion often increases mouth opening), the mask is expanded slightly downward beyond the chin. This aligns with widely used lower-face inpainting setups in Wav2Lip-style and diffusion-based lip-sync systems.
[…]
@@ 5 Experiments
Backbones. The visual encoder follows the Sapiens-0.3B configuration and is initialized from its pretrained weights; the decoder adopts the ViTMAE-B design with a cross-attention sublayer in every block to attend to the identity/ambient/vocal prompt tokens. The audio branch uses a pretrained wav2vec 2.0 base encoder kept frozen during pretraining. Pretraining. SyncLipMAE is trained on 128 GPUs with a per-GPU batch size of 48 (global $128{\times}48$) for 400k steps ( 15 days). Video dubbing fine-tuning. We initialize from WanVACE-1.3B, freeze Wan backbone blocks, and optimize only VACE blocks ( 0.7B trainable params). Training uses 128 GPUs, batch size 1 per GPU, for  30k steps ( 3 days). Each sample contains 17+1+117{+}1{+}1 frames: 17 target frames, one history frame (for temporal continuity), and one identity frame (appearance prior).
Pretraining data.
We pretrain SyncLipMAE on a mixture of talking-face corpora: Hallo3 Cui et al. (2025), CelebV-HQ Zhu et al. (2022), CelebV-Text Yu et al. (2023), MEAD Wang et al. (2020), VFHQ Xie et al. (2022), HDTF Zhang et al. (2021) and RAVDESS Livingstone & Russo (2018). Unless specified, audio is mono 16 kHz and videos are 25 fps. We adopt the official train–val–test split when available; otherwise, we randomly sample 100 instances each for validation and testing, and use the remainder for training.
@@
We evaluate on CelebV-HQ Zhu et al. (2022) using the emotion and action attribute sets, reporting video-level Accuracy and AUC (higher is better). Emotion is treated as a single-label, 8-way classification task, whereas action is a multi-label problem over 35 categories (one-vs-rest scoring aggregated at the video level). To match common practice, clip-based baselines (e.g., VideoMAE/MARLIN) adopt the $a{\times}b$ protocol—uniformly sample $a$ clips of $b$ frames and average their predictions following  Tong et al. (2022); frame-averaging models use $n$ uniformly sampled frames and average per-frame predictions Korbar et al. (2019). All methods use a single linear classifier on frozen features. Let $\{\mathbf{x}_{i}\}_{i=1}^{m}$ be the sampled clips (or frames), $f(\cdot)$ the frozen encoder, and $W$ the linear head; the video-level logits are
$\bar{\ell}(V)=\frac{1}{m}\sum_{i=1}^{m}W\,f(\mathbf{x}_{i}),\qquad\hat{y}(V)=\arg\max_{k}\mathrm{softmax}_{k}\!\big(\bar{\ell}(V)\big),$
with Accuracy/AUC computed from $\hat{y}(V)$ at the video level.
Results. Tab 3 summarises CelebV-HQ emotion and action. On emotion, SyncLipMAE clearly surpasses the best baseline with 16 frames and improves further at 80 frames, indicating benefits from longer temporal context. On action—where baselines are already strong—SyncLipMAE provides consistent but modest gains, with a slight boost when using 80 frames. Overall, SyncLipMAE delivers substantial improvements on emotion and steady gains on action.
@@
We evaluate SyncLipMAE on HDTF, VFHQ, and Hallo3, predicting text from silent videos. Following standard practice, accuracy is measured by Word Error Rate (WER), computed from word-level substitutions ($S$), deletions ($D$), and insertions ($I$) against the reference of length $N$:
@@
Results. Tab 5 reports WER on VFHQ, Hallo3, and HDTF. The visual-only head built on SyncLipMAE ’s tokens achieves the best WER on all three sets: $7.37$ on VFHQ, $5.99$ on Hallo3, and $5.78$ on HDTF, improving over Auto-AVSR by absolute $0.76$, $0.24$, and $0.13$ respectively, and substantially outperforming AV-HuBERT on each set. These gains indicate that SyncLipMAE ’s pretraining yields visual motion features that transfer effectively to sentence-level lip reading under the standard hybrid CTC/attention recipe.
@@
We evaluate video dubbing—re-synchronising the mouth region of a source video to a target speech while preserving identity, pose, and expressions—using a WanVACE-based model that supports both audio- and video-driven conditioning via a lightweight vocal-cue injection (details in section 4.4). We report the metrics in Tab 4. $\mathrm{Sync}_{\mathrm{conf}}$ (identical to LSE-C) is computed with the stable SyncNet evaluator from LatentSync Li et al. (2024), replacing the commonly used vanilla SyncNet whose scores are less reliable in our experiments Prajwal et al. (2020). Following SyncNet Chung & Zisserman (2016), for each short sliding window we form an audio–video distance curve over temporal offsets $\Delta\!\in\![-L,L]$:
@@
and define the per-window confidence as the median–minimum gap on this curve,

### v2

@@ 1 Introduction
Talking-face video underpins human–computer interaction, accessibility, content creation, and post-production. Practical systems must jointly reason about linguistic content, audio–visual timing, video editing to match a target speech track or reference motion, and facial expressions/actions. Classic synchronization pipelines such as SyncNet (syncnet) and subsequent Transformer-based variants (e.g., VocaLiST (vocalist)) focus on detecting in-/out-of-sync pairs or removing stream lags, while recent audio–visual pretraining methods (e.g., AV-HuBERT (avhubert)) have shown strong transfer to lip reading and AV-SR. Yet, these advances do not directly yield a token-level, stream-synchronised representation that cleanly separates identity, speech-synchronised mouth motion, and other facial dynamics—capabilities that are essential for both analysis and controllable generation.
Most prior work tackles one facet at a time. Visual or video masked modeling (MAE (mae)/VideoMAE (videomae)) excels at reconstruction but is modality-specific and does not enforce audio–visual alignment; contrastive AV learning improves correspondence but is typically framed as binary sync classification rather than a shared token space; and dubbing systems (e.g., Wav2Lip (wav2lip), LatentSync (latentsync), MuseTalk (musetalk)) optimise for generation quality and lip accuracy but are not designed to provide a unified interface that can be driven interchangeably by audio or reference motion within a single model. As a result, existing solutions often rely on wider temporal windows or task-specific heads, and they lack a factorised, alignment-aware representation that transfers seamlessly across synchronisation, understanding, and dubbing.
@@
Beyond analysis, SyncLipMAE enables unified visual dubbing: using a WanVACE-style (vace) DiT backbone for masked video inpainting, we inject either aligned audio tokens or vocal-motion tokens after patch embedding via a lightweight AudioPack, so the same model admits no-difference audio- or video-driven control. This preserves identity and pose while re-synchronising the mouth region, and it naturally benefits from the shared token space learned during pretraining.
@@ 2 Related Work
Audio–visual learning is often formulated as correspondence or temporal alignment. L3-Net l3net learns audio–visual correspondence from unlabelled video, while AVTS avts casts synchronisation as in-time vs. out-of-time discrimination within a clip. For faces, SyncNet syncnet introduces a two-stream embedding for lip–audio alignment and lag estimation, and Transformer-based models such as VocaLiST vocalist improve robustness across speech and singing with longer-range context.
@@
Contrastive pretraining aligns paired modalities at scale (e.g., CLIP (clip)), while masked autoencoders (MAE (mae), VideoMAE (videomae)) show that reconstructing heavily masked inputs yields strong visual and video representations. These two paradigms are now standard building blocks for multimodal pretraining.
@@
MARLIN (marlin) applies masked autoencoding to facial videos, using facial-region-guided masking to obtain a universal face encoder transferable to expression recognition, deepfake detection, and related tasks. Facial Region Awareness (FRA) fra further discovers facial regions via learned heatmaps and enforces consistency between global and local features, improving transfer across facial analysis benchmarks; subsequent variants adapt MAE-style pretraining to dynamic facial expression recognition under limited labels (mae_dfer). These methods focus on reconstruction-only objectives within the facial domain, whereas we couple MAE with audio–visual contrast to align speech-driven dynamics in talking-face video and explicitly separate identity and motion factors.
@@
End-to-end lipreading, large-scale audio–visual pretraining, and automatic recipe design have steadily improved VSR/AVSR. AV-HuBERT learns audio–visual speech representations from unlabelled video, and Auto-AVSR (autoavsr) automates model and loss selection for VSR/AVSR pipelines. ES3 (es3) decomposes audio–visual self-supervised learning into shared, modality-specific, and synergistic components and introduces an evolving Siamese training strategy, yielding strong low-resource VSR/AVSR. Our work is complementary: we also learn joint audio–visual speech–face representations, but optimise them for speech-driven facial dynamics and dubbing/synchronisation rather than word recognition.
@@
Audio-driven talking-head synthesis has progressed from GAN-based pipelines with explicit sync critics (e.g., Wav2Lip (wav2lip)) to latent/image-space diffusion (e.g., LatentSync (latentsync)) and efficient diffusion heads such as MuseTalk (musetalk). AV-HuBERT-based lip-sync experts avsr_expert instead treat a pretrained audio–visual speech model as a frozen teacher, using its features to define lip-sync losses and evaluation metrics. In parallel, foundation video generators and conditional DiT frameworks (DiT (dit), WAN (wan), VACE (vace)) provide scalable backbones and conditioning interfaces that we leverage in our dubbing setup. Unlike avsr_expert, we pretrain a dedicated face encoder jointly with audio rather than relying solely on a frozen AV speech expert.
@@ 3.1 Overview and Design Goals
SyncLipMAE aims to learn synchronization-aware talking-face representations from unlabeled audio–visual streams. As illustrated in figure 1(a), we combine MAE-style masked video modeling with a CLIP-style contrastive loss that pulls time-aligned vocal-motion and audio tokens together and pushes temporally misaligned pairs apart, yielding a shared token-level audio–visual space (§3.3). For each frame, a ViT encoder outputs three prompt tokens for identity $\mathbf{z}^{\mathrm{id}}$, vocal motion $\mathbf{z}^{\mathrm{voc}}$, and ambient motion $\mathbf{z}^{\mathrm{amb}}$, and the decoder reconstructs the frame via cross-attention to these prompts.
@@
Let $\mathbf{x}_{1:T}\in\mathbb{R}^{T\times C\times H\times W}$ be the video frames and $\mathbf{a}\in\mathbb{R}^{N}$ a mono waveform at sampling rate $f_{s}$. Each frame is patchified into $N_{p}$ patches and embedded as $D$-dimensional tokens. For MAE-style masking, let $\mathcal{V}_{t}\subset\{1,\dots,N_{p}\}$ be the visible indices at time $t$ and $\mathcal{M}_{t}$ the masked ones, with $|\mathcal{V}_{t}|=N_{\mathrm{vis}}$ and $\mathcal{V}_{t}\cup\mathcal{M}_{t}=\{1,\dots,N_{p}\}$; the visible patch tokens are $\mathbf{P}^{\mathrm{vis}}_{t}\in\mathbb{R}^{N_{\mathrm{vis}}\times D}$. A learnable mask token $\mathbf{e}^{\mathrm{mask}}\in\mathbb{R}^{D}$ is inserted at indices $\mathcal{M}_{t}$ to restore a length-$N_{p}$ sequence for decoding. Per frame, the encoder also outputs three prompt tokens $\mathbf{z}^{\mathrm{id}}_{t},\ \mathbf{z}^{\mathrm{amb}}_{t},\ \mathbf{z}^{\mathrm{voc}}_{t}\in\mathbb{R}^{D}$. For audio, we use a pretrained $L$-layer speech encoder kept frozen during pretraining; let $\{\mathbf{H}^{(\ell)}\}_{\ell=1}^{L}$ denote its hidden-state sequences. To align with the $T$ video frames, we resample $\mathbf{a}$ so that the encoder emits $T$ tokens per utterance at every layer, yielding $\mathbf{A}^{(\ell)}=\{\mathbf{a}^{(\ell)}_{t}\}_{t=1}^{T}$.
@@ 3.2 Model Design
Prompt Token Factorization Analysis.
To examine whether the three prompts specialize into identity, vocal motion, and ambient motion, we conduct two analyses. First, we reconstruct masked portraits while sourcing $\mathbf{z}^{\mathrm{id}}$, $\mathbf{z}^{\mathrm{amb}}$, and $\mathbf{z}^{\mathrm{voc}}$ from three frames with mismatched identity and expression (figure 3a) with the trained SyncLipMAE’s MAE Decoder. The reconstructions follow the ambient source in eye blinks (via $\mathbf{z}^{\mathrm{amb}}$) and the vocal source in mouth shape (via $\mathbf{z}^{\mathrm{voc}}$), while $\mathbf{z}^{\mathrm{id}}$ mainly controls static appearance. Second, we aggregate cross-attention weights from all decoder blocks and visualize per-prompt maps (figure 3b): attention from $\mathbf{z}^{\mathrm{voc}}$ concentrates on the lower face(mainly on the mouth), $\mathbf{z}^{\mathrm{amb}}$ spreads over the face with weak background response, and $\mathbf{z}^{\mathrm{id}}$ focuses on background and non-mouth/eye facial regions. These patterns provide qualitative evidence that the three prompts form a factorized representation over identity, vocal motion, and ambient motion.
@@ 3.3 Pretraining Objectives
Figure 2: (a) SyncLipMAE MAE reconstructions under prompt-source swapping. (b) Cross-attention maps for $\mathbf{z}^{\mathrm{id}}$, $\mathbf{z}^{\mathrm{amb}}$, and $\mathbf{z}^{\mathrm{voc}}$.
Figure 3: Our video dubbing architecture injects audio–visual synchronized feature to WanVACE DiT blocks.
@@ 4 Downstream Adaptation
Table 2: Used heads and prompt tokens.
Task Downstream head Tokens AV synchronisation Similarity (no head) $\mathbf{A}_{t}$, $\mathbf{z}^{\mathrm{voc}}_{t}$ Facial understanding Linear classifier $\mathbf{z}^{\mathrm{voc}}_{t}$, $\mathbf{z}^{\mathrm{amb}}_{t}$ Lip reading (VSR) Conformer+Transformer $\mathbf{z}^{\mathrm{voc}}_{t}$ Video dubbing WanVACE+AudioPack $\mathbf{A}_{t}$ / $\mathbf{z}^{\mathrm{voc}}_{t}$
We apply SyncLipMAE to four tasks (see Tab 2 for the downstream heads and tokens used in each case): (i) audio–visual stream synchronization — we estimate frame-level lag and synchrony in the frozen shared space by maximizing cosine similarity between aligned audio tokens $\mathbf{A}_{t}$ and vocal-motion tokens $\mathbf{z}^{\mathrm{voc}}_{t}$, without any additional head; (ii) facial understanding — we infer emotion and head/face actions by averaging the vocal- and ambient-motion tokens per frame and feeding a single linear classifier; (iii) visual speech recognition (VSR) — we infer text from the vocal-motion tokens using a Conformer encoder and Transformer decoder trained with a CTC/attention hybrid loss; (iv) visual dubbing — we edit a source face video to match target speech or a reference motion by conditioning a WanVACE DiT backbone via an AudioPack adapter that injects either audio tokens $\mathbf{A}_{t}$ or reference vocal-motion tokens $\mathbf{z}^{\mathrm{voc}}_{t}$, supporting audio- or video-driven control within a single model.
@@ 4.2 Facial understanding: emotion and action
We keep SyncLipMAE frozen and attach a minimal linear head. For each frame $t$, we form a motion descriptor by averaging the two motion prompts,
$\mathbf{u}_{t}=\tfrac{1}{2}\big(\mathbf{z}^{\mathrm{voc}}_{t}+\mathbf{z}^{\mathrm{amb}}_{t}\big),$
feed $\mathbf{u}_{t}$ into a single linear layer to obtain frame-level logits, and then average the logits over sampled frames before applying a softmax to obtain video-level predictions. The identity prompt is not used in this head, so the classifier operates purely on motion factors for emotion/action classification.
@@ 4.3 Lip Reading (VSR)
We adopt the visual speech recognition head of AutoAVSR (autoavsr) for lip reading and feed it with SyncLipMAE’s vocal-motion tokens. Concretely, let $\mathbf{V}=\{\mathbf{z}^{\mathrm{voc}}_{t}\}_{t=1}^{T^{\prime}}$ be the input sequence to a Conformer encoder (as in AutoAVSR), which produces hidden states $\mathbf{H}=\{\mathbf{h}_{t}\}_{t=1}^{T^{\prime}}$. A Transformer decoder then autoregressively predicts subword units with cross-attention over $\mathbf{H}$, following the standard encoder–decoder formulation.
@@
The total objective follows the standard hybrid CTC/attention form used in AutoAVSR,
@@ 4.4 Video dubbing
We cast dubbing as masked video inpainting on face crops within a VACE-style unified interface: task inputs (reference frames and a spatiotemporal mouth mask) are packed into a Video Condition Unit and processed by a Diffusion-Transformer backbone with Concept Decoupling and Context Adapters (vace). We instantiate this pipeline on WanVACE, keeping its architecture, losses, and training schedule unchanged; WanVACE natively accepts 1–3 reference images as a comma-separated list, which we use for identity control.
Control injection. We adopt an AudioPack placed after the VACE patch-embedding stage and before the first DiT VACE block, inspired by audio-conditioned injection in avatar generation (omniavatar). Per-frame control tokens—either audio tokens $\mathbf{A}_{t}$ from SyncLipMAE or reference vocal-motion tokens $\mathbf{z}^{\mathrm{voc}}_{t}$—are temporally encoded, projected to the DiT width, and added to the patch-embedded video token stream. No other WanVACE components are modified; switching $\mathbf{A}_{t}$ vs. $\mathbf{z}^{\mathrm{voc}}_{t}$ toggles audio-driven vs. video-driven control within the same model.
Long-form continuity via dual references. To support arbitrarily long, coherent dubbing, we provide two references: (i) an identity image (stable appearance prior) and (ii) the last frame of the previous segment (temporal bridge). During training we randomly drop the continuity reference with 10% probability so the model can also initialise the first segment without a history frame.
Masking strategy. We detect the facial region (chin to hairline) and mask its lower half as the editable area, expanding the mask slightly below the chin to accommodate jaw excursions. This follows common lower-face inpainting setups in Wav2Lip-style and diffusion-based lip-sync systems.
@@ 5 Experiments
Backbones. The visual encoder follows the Sapiens-0.3B (sapiens) configuration and is initialized from its pretrained weights; the decoder is a ViTMAE-B with a cross-attention sublayer in every block attending to the identity/ambient/vocal prompts. The audio branch uses a pretrained wav2vec 2.0 (wav2vec2) base encoder, kept frozen during SyncLipMAE training. SyncLipMAE training. SyncLipMAE is trained on 128 GPUs with batch size 48 per GPU (global $128{\times}48$) for 400k steps ($\sim$15 days), using loss weights $\lambda_{\mathrm{pix}}=\lambda_{\mathrm{cl}}=1$ and $\lambda_{\mathrm{cov}}=0.1$, temporal_neighbors =1=1 (3-frame positive window), and a self-similarity threshold of $0.9$. At each step we randomly sample a batch of frames from full videos, allowing consecutive or non-consecutive frames. Video dubbing fine-tuning. For dubbing we initialize WanVACE-1.3B, freeze all Wan backbone blocks, and optimize only the VACE blocks ($\sim$0.7B trainable parameters) for $\sim$30k steps ($\sim$3 days) on 128 GPUs with batch size 1 per GPU. Each sample contains 17+1+117{+}1{+}1 frames: 17 target frames, one history frame (temporal continuity), and one identity frame (appearance prior).
SyncLipMAE training data.
We pretrain SyncLipMAE on Hallo3 hallo3, CelebV-HQ celebvhq, CelebV-Text celebvtext, MEAD mead, VFHQ vfhq, HDTF hdtf, and ESS ravdess. Audio is converted to mono 16 kHz and videos are resampled to 25 fps. For VFHQ and MEAD we follow the official evaluation protocols, using the released test partitions for evaluation and the remaining clips for pretraining, and additionally sampling 100 training clips for validation. For Hallo3, CelebV-HQ, CelebV-Text, HDTF, and RAVDESS, which lack standardized splits in our setting, we randomly sample 100 clips for validation and 100 clips for testing, and use the rest for pretraining. All videos are decoded, uniformly resized, and center-cropped to $512{\times}512$, ensuring that the full face remains.
@@
Method VFHQ Hallo3 $\mathrm{Sync}_{\mathrm{conf}}\,\uparrow$ FID $\downarrow$ FVD $\downarrow$ $\mathrm{Sync}_{\mathrm{conf}}\,\uparrow$ FID $\downarrow$ FVD $\downarrow$ MuseTalk musetalk 0.6820 24.44 175.16 0.6904 13.80 86.46 LatentSync-1.5 latentsync 0.7681 17.92 131.04 0.7725 10.59 78.89 A-SyncLipMAE (ours) 0.7830 16.46 100.44 0.8027 10.04 74.01 V-SyncLipMAE (ours) 17.22 109.91 10.82 77.57
@@
We evaluate on CelebV-HQ (celebvhq) using the emotion and action attribute sets, reporting video-level Accuracy and AUC. All classifiers are trained on the CelebV-HQ training split and evaluated on the test split. Emotion is an 8-way single-label task, whereas action is multi-label over 35 categories (one-vs-rest at the video level). Following common practice (videomae; scsampler), clip-based baselines (e.g., VideoMAE/MARLIN) use an $a{\times}b$ protocol—uniformly sampling $a$ clips of $b$ frames and averaging clip predictions; frame-averaging models sample $n$ frames and average per-frame predictions. All methods attach a single linear classifier to frozen features: given sampled clips or frames $\{\mathbf{x}_{i}\}_{i=1}^{m}$, encoder $f$, and head $W$,
$\bar{\ell}(V)=\frac{1}{m}\sum_{i=1}^{m}W\,f(\mathbf{x}_{i}),$
with video-level predictions and Accuracy/AUC computed from $\bar{\ell}(V)$.
[…]
Results. Tab 4 summarizes CelebV-HQ emotion and action. On emotion, SyncLipMAE clearly surpasses the best baseline with 16 frames and improves further at 80 frames, indicating benefits from longer temporal context. On action—where baselines are already strong—SyncLipMAE provides consistent but modest gains, with a slight boost when using 80 frames. Overall, SyncLipMAE delivers substantial improvements on emotion and steady gains on action.
@@
We evaluate SyncLipMAE on HDTF, VFHQ, and Hallo3 for visual-only VSR. A single VSR head is trained jointly on the training splits of all three datasets, using speech transcripts obtained by running the Whisper Turbo model (whisper) on the original audio, and evaluated on the corresponding test splits after filtering out non-English utterances. Following standard practice, we report Word Error Rate (WER), computed from word-level substitutions ($S$), deletions ($D$), and insertions ($I$) against a reference of length $N$:
@@
Results. Tab 6 reports WER on VFHQ, Hallo3, and HDTF. For AV-HuBERT, we use the official Large + Self-Training checkpoint pretrained on LRS3 + VoxCeleb2 (En) and finetuned on LRS3-433h. The visual-only head built on SyncLipMAE ’s tokens achieves the lowest WER on all three sets: $7.37$ on VFHQ, $5.99$ on Hallo3, and $5.78$ on HDTF, improving over Auto-AVSR by absolute $0.76$, $0.24$, and $0.13$, respectively, and outperforming AV-HuBERT on each set. These results indicate that SyncLipMAE ’s pretraining yields visual motion features that transfer effectively to sentence-level lip reading under the standard hybrid CTC/attention recipe.
@@
We train a WanVACE-based video dubbing model on the same mixture of talking-face corpora used for SyncLipMAE pretraining and evaluate it on the HDTF, Hallo3, and VFHQ test splits, re-synchronising the mouth region of a source video to a target speech while preserving identity, pose, and expressions. The model follows a VACE-style unified interface and supports both audio- and video-driven conditioning via a lightweight vocal-cue injection (details in section 4.4). We report the metrics in Tab 5. $\mathrm{Sync}_{\mathrm{conf}}$. Identical to LSE-C, it is computed with the stable SyncNet evaluator from LatentSync (latentsync), instead of the vanilla SyncNet whose scores we found less reliable (syncnet). Following SyncNet, for each sliding window we form an audio–video distance curve over temporal offsets $\Delta\!\in[-L,L]$:
@@
and define the per-window confidence as the median–minimum gap,